### v2

# Number it: Temporal Grounding Videos like Flipping Manga

Abstract
Video Large Language Models (Vid-LLMs) have made remarkable advancements in comprehending video content for QA dialogue. However, they struggle to extend this visual understanding to tasks requiring precise temporal localization, known as Video Temporal Grounding (VTG). To address this gap, we introduce Number-Prompt (NumPro), a novel method that empowers Vid-LLMs to bridge visual comprehension with temporal grounding by adding unique numerical identifiers to each video frame. Treating a video as a sequence of numbered frame images, NumPro transforms VTG into an intuitive process: flipping through manga panels in sequence. This allows Vid-LLMs to “read” event timelines, accurately linking visual content with corresponding temporal information. Our experiments demonstrate that NumPro significantly boosts VTG performance of top-tier Vid-LLMs without additional computational cost. Furthermore, fine-tuning on a NumPro-enhanced dataset defines a new state-of-the-art for VTG, surpassing previous top-performing methods by up to 6.9% in mIoU for moment retrieval and 8.5% in mAP for highlight detection. The code will be available at https://github.com/yongliang-wu/NumPro.
Figure 1: Effectiveness of Adding Frame Numbers for Temporal Grounding: (a) Without numbered images or frames, both humans and Vid-LLMs struggle to locate specific timestamps accurately. (b) Once numbered, grounding temporal cues becomes as intuitive as flipping manga, where timestamps are accessible at a glance.
1 Introduction
Imagine you are watching a cooking video, and trying to locate the exact moment when the chef stirs in the spices. While recognizing such actions is feasible, translating that visual information into precise timing, i.e., a specific second or frame number, is surprisingly difficult. This challenge is central to the field of Video Temporal Grounding (VTG) [17, 4, 43, 47, 22]. In the realm of Video Large Language Models (Vid-LLMs) [55, 36, 71, 29], the integration of VTG allows for fine-grained visual and temporal understanding and reasoning of videos, which is pivotal for developing end-to-end video dialogue systems.
Despite advances of Vid-LLMs, endowing these models with effective VTG abilities presents a unique challenge: enhancing the model’s visual recognition of an event within a video does not inherently enable it to describe when the event begins and ends using language [22, 47]. For instance, advanced Vid-LLMs like Qwen2-VL [55], while excelling at video comprehension, can struggle with grounding specific events in time. When asked, e.g., to locate “when does the woman eat food” in a 10-frame video, the model can hallucinate an illogical answer like “from frame 000 to 580.” [^1] See more cases in Appendix 7.1. This limitation arises because these models are primarily trained to align visual content with language descriptions (what happens) while lacking mechanisms to directly interpret the temporal boundaries (when does it happen).
This gap in powerful Vid-LLMs leads us to think: How can we empower Vid-LLMs to extract temporal cues directly through visual recognition? A familiar human experience – flipping manga – provides an intuitive solution. When flipping manga, each numbered panel guides readers to follow the sequence of the narrative, linking visual content with a clearly defined timeline. Inspired by this, we introduce Number-Prompt (NumPro), which places unique numerical identifiers on each video frame, similar to manga panel numbers.
Current Vid-LLMs process videos by considering them as a sequence of frame images. With NumPro, VTG is as intuitive as flipping manga. As shown in Figure 1, NumPro augments each frame with a unique numerical identifier denoting its position in the temporal sequence. Given a language query targeting an event, Vid-LLMs retrieve relevant visual features of video frames and associate with the frame numbers overlaid. These numerical identifiers are then directly translated into textual outputs. In practice, we strategically position frame numbers in the bottom-right corner, using a medium font size and distinct color. This design ensures numbers visibility without obstructing essential visual content. Overall, NumPro allows Vid-LLMs to “read” the video timeline, effectively converting visual recognition into a temporal narrative in language.
NumPro’s elegance lies in its simplicity: by subtly adding frame numbers as temporal markers into video frames, we enable Vid-LLMs to naturally correlate each frame to its temporal location in the video sequence. Unlike previous approaches [47, 22, 18, 56, 33, 45, 23], NumPro does not introduce additional tokens or modify model vocabulary to provide temporal cues, thus avoiding additional learning complexities and maintaining strong transferability across various tasks and datasets. Temporal grounding, therefore, becomes an accessible, “free-lunch” enhancement for Vid-LLMs already proficient in understanding video content. Additionally, fine-tuning on a specially curated NumPro-enhanced VTG dataset (NumPro-FT) further advances state-of-the-art performance.
Our contributions can be summarized as follows:
  • We introduce NumPro, a novel approach that enhances Video Temporal Grounding (VTG) capabilities of Video Large Language Models (Vid-LLMs) by overlaying frame numbers onto video frames, making temporal grounding as intuitive as following numbered panels in flipping manga.
  • Through an experimental study, we find a suitable NumPro design (font size, color, and position) that ensures high detectability by the model while minimally interfering with the original video content.
  • We thoroughly evaluate NumPro on standard VTG benchmarks and metrics in both training-free and fine-tuned scenarios, demonstrating its effectiveness across various models and datasets.
2 Related Work
Video Temporal Grounding with Vid-LLMs. Video Temporal Grounding (VTG) [43, 54, 72, 32] focuses on the precise identification of event timestamps within videos, covering tasks such as moment retrieval [17, 16, 3, 68, 53, 4, 7, 63, 14, 62, 37], dense captioning [26, 24, 64, 53, 4, 8], and highlight detection [28, 47]. For current Video Large Language Models (Vid-LLMs) [55, 71, 29], which leverage powerful LLMs [1] for cross-modal understanding and video-based reasoning, VTG is crucial for achieving fine-grained temporal and visual comprehension, enabling end-to-end video dialogue systems with integrated temporal reasoning [22, 47, 56]. To achieve this, some methods rely on refined instruction datasets with temporal information (timestamps or frame numbers) for model fine-tuning [22, 33], while others concatenate additional textual temporal timestamps tokens with visual inputs [20, 50] or introduce specific temporal embeddings [47, 18]. Additional strategies model video structure [19, 23, 56] to better segment or organize videos into parts suitable for VTG. However, these approaches often require extensive retraining or specialized model adaptations, limiting their flexibility and transferability. In contrast, our NumPro aims to improve VTG for existing Vid-LLMs without additional training costs or architectural modifications.
Visual Prompt in VLMs. Visual prompts [58], taking various forms such as circles [5, 65], bounding boxes [13, 10, 38] and semantic masks [67, 40], enhance vision-language models (VLMs) [39, 66, 70, 60, 51, 57] to focus on and reason about specific visual regions and reduce the occurrence of hallucination [5]. For CLIP [46], a simple red circle [49] or colored region [67] can effectively guide model attention. Multi-modal large language models (MLLMs) [31, 2, 6] are also sensitive to specific visual prompts [5]. For example, ViP-LLaVA [5] and SoM [65] prompt MLLMs to answer about specific image regions with graphic shapes or numeric tags. CoLLaVO [27] and DOrA [59] utilize pixel-level prompts in images or videos to enhance the semantic localization capability of MLLMs. Additionally, toolchain [61, 73, 52, 48] approaches aggregate various visual prompts into multi-step reasoning paradigm to support reasoning complex tasks. While prior works focus on enhancing the region-based visual understanding of VLMs with visual prompts, our NumPro is the first to employ simple numerical tags as visual prompts within video frames to improve the temporal grounding capability.
Figure 2: Attention Analysis between Video Frames and Event Query. Although the model accurately attends to regions of interest related to the query, it struggles to generate precise temporal boundaries in its response.
Figure 3: Framework of Our Approach in Two Settings: (1) Training-free VTG with NumPro, where frame numbers are directly added to video frames, enabling Vid-LLMs to locate events temporally without additional training, and (2) Fine-tuned VTG with NumPro-FT, which further improves VTG performance by fine-tuning Vid-LLMs on a dataset NumPro-enhanced with no architectural modifications.
3 Number-Prompt Approach
Our Number-Prompt (NumPro) approach provides a simple yet effective solution to enhance Video Temporal Grounding (VTG) capabilities of existing Video Large Language Models (Vid-LLMs) in both training-free and fine-tuned settings, as shown in Figure 3. Section 3.1 presents an attention analysis based on Qwen2-VL [55] to highlight the challenge of aligning visual features with textual temporal boundaries. Section 3.2 describes the construction of NumPro and the fine-tuning process of Vid-LLMs on a NumPro-augmented VTG dataset, referred to as NumPro-FT. Finally, Section 3.3 details the design optimization of NumPro for maximizing its effectiveness.
3.1 Attention Analysis
Current Vid-LLMs process videos as a sequence of frames. Visual representations of the video can be taken as the concatenated representations from each individual frame, aggregating the information from discrete frames into a comprehensive video level. This allows Vid-LLMs to understand videos by aligning visual representations of frame images with the textual representations of language queries.
To explore the challenge in video temporal grounding (VTG), we analyze the attention map between representations of the frame image tokens and the query language tokens, and then we assess the temporal description of relevant video frames. Using Qwen2-VL-7B [55] as a case study, we highlight the challenge of VTG for Vid-LLMs: while Vid-LLMs can understand what event is happening within a video, they struggle to translate this understanding into a textual description that describes when the event begins and ends.
Specifically, we take a video and a language query as input, and extract the attention scores from the final multi-head self-attention layer of Qwen2-VL-7B [55]. For each frame within the video sequence, we aggregate the attention scores from all the visual tokens corresponding to that frame across all attention heads.
As illustrated in Figure 2, the attention map reveals a strong correlation between the text query of an event and targeted video segments. It indicates that Qwen2-VL-7B can effectively focus on query-relevant frames, which is consistent with the model’s strong performance in other content-related video understanding tasks [15, 30]. However, the model struggles to verbalize the correct temporal boundaries, and generates surprising hallucinations such as “from 200 to 599.” [^2] See more cases in Appendix 8. This observation underscores the need for mechanisms that can bridge the gap between spatial feature alignment and temporal reasoning with Vid-LLMs, which we aim to address with our NumPro method.
3.2 NumPro and NumPro-FT
Our approach, Number-Prompt (NumPro), empowers Vid-LLMs to directly associate specific visual content with its temporal information, turning temporal localization into a visual alignment task. As shown in Figure 3, NumPro operates in both training-free and fine-tuned scenarios.
In the training-free setting, each video frame is marked with its corresponding frame number. By utilizing the built-in Optical Character Recognition (OCR) capabilities of Vid-LLMs, we enable them to “read” the timeline through the frame numbers associated with visual content. To clarify the purpose of the added numbers to Vid-LLMs, we prepend a simple instruction to each event query: “The red numbers on each frame represent the frame number.” This approach allows Vid-LLMs to accurately identify frame-level boundaries by directly linking the frame numbers to language queries.
For improved performance, NumPro-FT fine-tunes Vid-LLMs on a NumPro-augmented dataset. This stage aligns frame numbers with temporal spans within the training data, embedding temporal grounding capabilities into the model’s learned representations. During fine-tuning, we freeze the visual encoder and only fine-tune the visual projector and LLM components. To reduce parameter count and training overhead, we apply Low-Rank Adaptation (LoRA) [21] to adjust the LLM. Our training objective is to maximize the likelihood of generating the correct answer tokens $\mathbf{A}$ via auto-regressive language modeling:
$P(\mathbf{A}\mid V,T_{\text{instruct}})=\prod_{j=1}^{L}P_{\theta}(A_{j}\mid V,X_{\text{instruct}},\mathbf{A}_{<j})$ (1)
Here, $V$ represents the input video, $\theta$ denotes the trainable parameters, $T_{\text{instruct}}$ is the text instruction, $L$ is the length of the answer sequence $\mathbf{A}$, and $\mathbf{A}_{<j}$ includes all preceding answer tokens before the current token $A_{j}$.
3.3 Design of Numerical Prompt
An effective NumPro design must ensure: (1) numbers are easily recognized by the model, and (2) minimal interference with visual content. Previous research [5] indicates that the appearance and placement of visual prompts can influence model attention. Given that all Vid-LLMs operate at a fixed resolution of $336\times 336$, we optimize NumPro by assessing three factors: font size, color, and placement position of the frame number.
To determine an effective NumPro design, we use two primary metrics: Number Accuracy, assessing how well the model identifies overlaid numbers, and Caption Accuracy, measuring how accurately the original caption aligns with frame content after adding numbers. Balancing these two metrics allows us to select NumPro configurations where the numbers are clearly recognizable without disrupting the main video content.
To make the design choices robust across various models and datasets, we employ CLIP-based experiments on a subset of MSCOCO dataset [35] to calculate Number Accuracy and Caption Accuracy separately. We use the CLIP ViT-B/32 [46, 11] model to generate visual and textual representations, as many Vid-LLMs utilize CLIP-style vision encoders [55, 12], allowing our findings to generalize well across Vid-LLMs. COCO image-caption pairs serve as proxies for video frames, avoiding the high costs and limited scalability of direct VTG testing. Specifically, we randomly select 1,000 distinct image-caption pairs from MSCOCO [35] and overlay numbers ranging from “0” to “99” onto the image in various configurations.
As shown in Figure 4, we first obtain representations from CLIP [46] vision and text encoders and compute intermediate similarity scores (i.e., Number and Caption Similarity) between them. Using the added numbers and original captions as ground truth, we select the text numbers and captions with the highest similarity scores as predictions to calculate Number and Caption Accuracy. Configurations balancing these accuracies are optimal for NumPro design.
Figure 4: Illustration of Our NumPro Design Algorithm.We overlay different numbers onto COCO images and obtain visual and textual representations using CLIP encoders. For each configuration, we calculate Number/Caption Similarity and derive Number/Caption Accuracy, enabling us to identify the optimal NumPro design that balances recognizability and minimal disruption to the visual content.
Figure 5: The Impact of Different Number-Prompt Designs. We categorize the design into three dimensions: font size, position, and color. BL stands for Bottom Left, BR for Bottom Right, TL for Top Left, TR for Top Right, and C for Center.
As shown in Figure 5, our findings indicate that increasing the font size improves number accuracy but reduces caption accuracy, suggesting that a moderate font size (40 or 60) is optimal. For color selection, caption accuracy remains relatively stable across different colors. Red shows the best performance for number accuracy, while black was the least effective. This finding is also consistent with previous works [49, 5]. Additionally, positioning the text in the center of the image significantly reduced caption accuracy due to overlaps with key visual elements, while placing the numbers in the bottom-right corner provides the best balance between caption and number accuracy. Finally, we select a font size of 40, the color red, and the bottom-right position for our final NumPro design. This design search allows NumPro to better harness the inherent OCR and visual-language alignment capabilities of Vid-LLMs to empower video temporal grounding.
In practice, CLIP-based designs provide approximate rather than definitive guidance, further testing on Vid-LLMs with a VTG dataset may yield additional model-specific insights. In Sec 4.3, consistent results further validate the effectiveness of our design.
4 Experiments
Table 1: Comparison of performance on the video temporal grounding task with previous state-of-the-art methods. NumPro refers to the use of number prompts for augmentation during inference, while NumPro-FT indicates fine-tuning with the number prompt augmentation instruction dataset. The best results are highlighted in bold, and the second-best are underlined.
Model Charades-STA ActivityNet QVHighlights
R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 mIoU mAP HIT@1
VTG-Tuned Vid-LLMs
GroundingGPT [33] - 29.6 11.9 - - - - - - -
LITA [23] - - - - - 25.9 - 28.6 - -
VTG-LLM [18] 52.0 33.8 15.7 - - - - - 16.5 33.5
TimeChat [47] 47.7 22.9 12.5 30.6 30.2 16.9 8.2 21.8 14.5 23.9
VTimeLLM [22] 51.0 27.5 11.4 31.2 44.0 27.8 14.3 30.4 - -
Momentor [45] 42.9 23.0 12.4 29.3 42.6 26.6 11.6 28.5 7.6 -
HawkEye [56] 50.6 31.4 14.5 33.7 49.1 29.3 10.7 32.7 - -
General Vid-LLMs
GPT-4o [44] 55.0 32.0 11.5 35.4 33.3 21.2 10.4 23.7 39.5 68.7
        +NumPro 57.1 35.5 13.5 37.6 45.5 30.8 18.4 33.6 40.5 70.7
Qwen2-VL-7B [55] 8.7 5.4 2.4 7.9 17.0 9.4 3.9 12.5 21.5 42.2
        +NumPro 60.7 36.8 15.9 38.5 44.2 26.4 14.4 31.3 23.6 43.4
LongVA-7B-DPO [69] 22.6 10.1 2.2 14.6 11.8 5.3 1.9 8.2 14.2 20.4
        +NumPro 27.2 10.3 2.9 18.9 20.1 10.8 5.4 15.2 15.3 24.3
        +NumPro-FT 63.8 42.0 20.6 41.4 55.6 37.5 20.6 38.8 25.0 37.2
We evaluate our model on two Video Temporal Grounding (VTG) tasks: Moment Retrieval [4, 17] and Highlight Detection [28]. Moment Retrieval, given a language query describing an event, identifies the specific start and end video frames of the event. We utilize Charades-STA [17] and ActivityNet [4] as evaluation datasets, following previous works [22, 47, 45, 56]. Evaluation metrics include the mean Intersection over Union (mIoU) and recall@1 at various IoU thresholds $m$ (R@$m$), where $m$ is set to {0.3, 0.5, 0.7} following previous work [22, 47]. For Highlight Detection, which aims to locate and rank video frames based on their relevance to the language query, we use QVHighlights [28] for evaluation. Evaluation metrics include mean Average Precision (mAP) and HIT@1 (the hit ratio of the highest-scored clip), as in  [47, 28, 18]. Please see Appendix 11 for more task examples.
4.1 Implementation Details for NumPro-FT
Dataset Preparation. Our temporal grounding dataset consists of 70k question-answer pairs from DiDeMo [42] and ActivityNet Caption [4] datasets. Additionally, we incorporate data from Stage 2 and Stage 3 of VTimeLLM dataset [22]. After filtering out invalid videos, we obtain a comprehensive instruction dataset totaling 220k samples. Each video in our dataset is augmented with our NumPro method by overlaying frame numbers directly onto the video frames. The question-answer pairs follow a consistent template: questions are formatted as “During which frames can we see {query}?” and answers are formatted as “From $x$ to $y$”, where $x$ and $y$ denote the start and end frame numbers of the query event.
Training Details. We utilize the LongVA-7B-DPO [69] as our base model, taking into account its uncomplicated design and its extensive capacity to handle context length. Additionally, it has not been trained on any video data. The model is trained for 3 epochs over our curated dataset with a total batch size of 128. We use the AdamW optimizer [25] with cosine learning rate decay. The learning rate is set to 1e-4, and the warm-up ratio is 0.05. The LLM component utilizes LoRA with parameters $r=64$ and $\alpha=128$. All experiments are conducted on 8 H100 GPUs.
4.2 Main Results
4.2.1 Comparison with State-of-the-Art Methods
Table 1 presents a comparative analysis of Vid-LLMs enhanced with our NumPro/NumPro-FT against existing state-of-the-art (SOTA) methods on Moment Retrieval and Highlight Detection tasks.
Moment Retrieval: Applying training-free NumPro enables Vid-LLMs to approach or exceed previous SOTA performance, benefiting both closed-source and open-source Vid-LLMs. GPT-4o [44] already exhibits strong moment retrieval capabilities, and our NumPro further enhances the performance. In particular, NumPro achieves a 9.9% increase in mIoU on ActivityNet, surpassing the previous SOTA by 0.9%. Qwen2-VL-7B performs poorly initially and also sees a significant improvement with NumPro, averaging a 24.7% increase in mIoU across datasets.
Moreover, starting from a relatively low baseline on LongVA-7B-DPO [69], our fine-tuning approach, NumPro-FT, establishes new SOTA across all metrics. On Charades-STA, it surpasses previous SOTA by 11.8%, 8.2%, 4.9%, 7.7% (R@0.3, R@0.5, R@0.7, mIoU), and on ActivityNet, it surpasses previous SOTA by 6.5%, 8.2%, 6.3%, 6.1% (R@0.3, R@0.5, R@0.7, mIoU). These results demonstrate that NumPro and NumPro-FT can utilize the superior video understanding abilities of existing Vid-LLMs and significantly enhance their moment retrieval capabilities.
Highlight Detection: In this task, models like GPT-4o [44] and Qwen2-VL have already achieved state-of-the-art (SOTA) performance. However, our NumPro approach consistently enhances their performance, with an average increase of 1.55% in mean Average Precision (mAP) and 1.6% in the hit ratio of the highest-scored clip (HIT@1). Additionally, applying NumPro-FT enables LongVA-7B-DPO to surpass existing SOTA by a large margin (+8.5% in mAP and +3.7% in HIT@1). These findings suggest that NumPro and NumPro-FT, which can be easily appended to current Vid-LLMs, hold substantial potential for further advancing temporal reasoning capabilities.
4.2.2 Effectiveness of NumPro across Vid-LLMs
Table 2: Performance of Applying NumPro to Various Vid-LLMs and Ablation Results on NumPro-FT.
Model Charades-STA ActivityNet
R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 mIoU
LLaVA-OneVision-7B [29] 22.3 7.9 2.1 15.9 7.1 3.1 1.1 6.1
        +NumPro $42.9_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+20.6)}}}$ $19.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+11.5)}}}$ $6.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+4.5)}}}$ $28.1_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+12.2)}}}$ $14.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+7.3)}}}$ $7.9_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+4.8)}}}$ $3.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+2.7)}}}$ $11.3_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+5.2)}}}$
LLaVA-Video-7B [71] 11.8 2.7 0.1 9.8 7.4 3.1 1.2 6.2
        +NumPro $56.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+44.8)}}}$ $25.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+22.9)}}}$ $8.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+8.5)}}}$ $34.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+24.8)}}}$ $25.2_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+17.8)}}}$ $15.2_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+12.1)}}}$ $8.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+7.2)}}}$ $18.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+12.4)}}}$
Qwen2-VL-72B [55] 0.0 0.0 0.0 0.2 1.0 0.6 0.3 1.0
        +NumPro $25.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+25.8)}}}$ $9.9_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+9.9)}}}$ $3.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+3.0)}}}$ $17.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+17.2)}}}$ $35.5_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+34.5)}}}$ $21.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+20.8)}}}$ $11.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+10.7)}}}$ $25.5_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+24.5)}}}$
LongVA-7B-DPO [69] 22.6 10.1 2.2 14.6 11.8 5.3 1.9 8.2
        +FT 62.0 41.6 19.9 40.2 41.8 25.7 13.7 29.0
        +NumPro-FT $63.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+41.2)}}}$ $42.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+31.9)}}}$ $20.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+18.4)}}}$ $41.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+26.8)}}}$ $55.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+43.8)}}}$ $37.5_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+32.2)}}}$ $20.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+18.7)}}}$ $38.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+30.6)}}}$
Beyond surpassing SOTA, Table 2 demonstrates the broad applicability and scalability of NumPro across various Vid-LLMs in Video Temporal Grounding. We apply NumPro to additional Vid-LLMs, including LLaVA-Video-7B [71], LLaVA-OneVision-7B [29], and Qwen2-VL-72B [55], and observe notable performance improvements, with average mIoU gains reaching up to 18.1% on Charades and 14.0% on ActivityNet. Moreover, we conduct fine-tuning experiments with and without NumPro-augmented data (indicated as +FT in Table 2). Results show that NumPro-FT consistently outperforms conventional fine-tuning, particularly on longer video datasets like ActivityNet, where it achieves a substantial 9.8% gain in mIoU. Additional studies on NumPro’s effectiveness for QVHighlights are provided in Appendix 9. Those observations underscore the effectiveness of NumPro across models and highlight its superior impact when combined with fine-tuning.
4.2.3 Qualitative Results
In Figure 6, we compare our method with SOTA methods, TimeChat [47] and VTimeLLM [22], through two visualization cases from our ActivityNet dataset. The first example features minimal scene changes between video frames. TimeChat predicts an early start, while VTimeLLM fails to capture the full event duration. In contrast, our method precisely captures the correct event boundaries. The second case involves a shorter event duration and frequent scene changes. TimeChat completely misses the event, and VTimeLLM overestimates the event duration by including irrelevant segments. Our approach, again, precisely delineates the event boundaries. These qualitative examples underscore the robustness and precision of our method in scenarios that are especially challenging for other SOTA methods. We provide additional cases on moment retrieval and highlight detection in Appendix 11.
Figure 6: Qualitative Comparison with State-of-the-Art. Our LongVA-7B-DPO model, fine-tuned with NumPro-FT, outperforms TimeChat [47] and VTimeLLM [22] on ActivityNet by accurately identifying event boundaries in challenging scenes.
Table 3: Ablation study on various NumPro designs. We divide the designs into three dimensions: font size, color, and position.
Size Color Position Charades-STA
R@0.3 R@0.5 R@0.7 mIoU
40 Red Top Left 56.7 32.9 13.8 35.8
40 Red Top Right 58.2 34.0 13.0 36.8
40 Red Center 53.7 29.5 10.4 34.1
40 Red Bottom Left 61.6 37.8 15.9 39.3
40 Red Bottom Right 60.7 36.8 15.9 38.5
20 Red Bottom Right 53.6 34.0 14.0 34.6
40 Red Bottom Right 60.7 36.8 15.9 38.5
60 Red Bottom Right 58.0 34.5 14.1 37.1
80 Red Bottom Right 58.0 33.9 13.7 36.9
40 Red Bottom Right 60.7 36.8 15.9 38.5
40 Blue Bottom Right 57.8 34.2 14.6 36.6
40 Black Bottom Right 56.6 36.0 15.9 36.6
40 Green Bottom Right 56.0 33.8 14.5 36.0
4.3 Validation of NumPro Design
Following our heuristic design process in Sec 3.3, we validate its effectiveness in temporal grounding tasks to confirm that these design choices generalize beyond the COCO dataset. We conduct moment retrieval experiments on Charades-STA [17] with Qwen2-VL-7B [55] in a training-free setting. As shown in Table 3, the results align closely with our initial observations from COCO dataset, confirming the effectiveness of our design choices in VTG tasks. Specifically, (1) Position: Consistent with our CLIP-based findings, placing the text in the center has the largest impact on performance due to overlaps, while our choice of the bottom-right performs comparably to the best position; (2) Font Size: Both very large and very small fonts yield suboptimal results, supporting our balanced selection; (3) The performance on VTG is sensitive to number color, yet the red color consistently delivers the best performance, which may attribute to its high contrast against typical backgrounds [49]. Overall, the alignment between the CLIP-based design choices and the VTG results shows the validity and robustness of our NumPro design. Please refer to Appendix 10 for the ablation results of NumPro-FT.
Figure 7: Performance Comparison of Sampling Strategies for NumPro. We compare the effects of NumPro with different sampling ratios and sampling methods (random vs. uniform), as tested on the Charades-STA [17] using the Qwen2-VL-7B [55] model.
4.4 Investigation on the Sampling of NumPro
Typically, we augment every frame in a video with NumPro. In this section, we evaluate the impact of varying the sampling ratio and sampling method (randomly or uniformly) when selecting a subset of frames from the video to augment NumPro. As depicted in Figure 7, performance increases with more labeled frames, with uniform sampling generally maintaining higher accuracy. Notably, labeling just 20% of the frames provides a substantial performance boost and uniform sampling of 80% of the frames surpasses previous state-of-the-art, underscoring the robustness of our NumPro approach.
4.5 Influence on General Video-QA
To explore the broader applicability of NumPro, we integrate it into general video-QA tasks, using VideoInstruct [41] as our benchmark. As detailed in Table 4, the incorporation of NumPro minimally affects general comprehension metrics, with a slight decrease in Distraction Overlap (DO, -0.02) and an enhancement in Temporal Understanding (TU, +0.1). This indicates that Vid-LLMs equipped with NumPro maintain robust performance in general video-QA while excelling in precise video temporal grounding (VTG) tasks. This dual capability allows us to harness a powerful Vid-LLM for end-to-end video understanding that can flexibly adapt to both general and temporally nuanced questions within conversational AI systems.
Table 4: The influence of applying NumPro to general video-QA. CI stands for correctness of information, DO stands for detail orientation, CU stands for contextual understanding, TU stands for temporal understanding, and CO stands for consistency.
Model CI DO CU TU CO
Qwen2-VL 3.10 2.57 3.46 2.47 3.30
        +NumPro 3.10 2.55 3.46 2.57 3.30
5 Conclusion
In this paper, we propose Number-Prompt (NumPro), a simple yet efficient visual prompt designed to enhance the video temporal grounding (VTG) capabilities of Video Large Language Models (Vid-LLMs) with no effort. By overlaying frame numbers onto video content, NumPro leverages the inherent Optical Character Recognition (OCR) and visual-language alignment capabilities of Vid-LLMs, allowing them to accurately map events to specific temporal boundaries. Through systematic design informed by COCO-based heuristics and validated across VTG benchmarks, we demonstrated that NumPro effectively supports fine-grained temporal understanding while preserving general video comprehension. Through extensive evaluations, we demonstrated that NumPro consistently achieves state-of-the-art performance in both training-free and fine-tuned settings, enabling adaptable integration into both closed-source and open-source Vid-LLMs. NumPro-FT further refines temporal grounding performance, establishing new SOTA across VTG tasks. Besides, the minimal impact on general video-QA shows that NumPro can augment VTG while maintaining robust video understanding.
Supplementary Material
6 Experimental Details
6.1 Moment Retrieval
In the training-free (NumPro) setting, we extract frames from videos at 1 FPS, with each frame resized to a resolution of $336\times 336$. In the fine-tuned (NumPro-FT) setting, frames are extracted at 0.5 FPS during both the training and inference phases due to GPU memory constraints.
6.2 Highlight Detection
In both the training-free (NumPro) and fine-tuned (NumPro-FT) settings, we extract frames from videos at 0.5 FPS because the saliency score is labeled every 2 seconds. Each frame is resized to a resolution of $336\times 336$.
7 Hallucination in Vid-LLMs for VTG
7.1 General Vid-LLMs
Qwen2-VL-7B. Figure 8(a) shows the results of Qwen2-VL-7B [55] suffer from severe hallucinations. For instance, the model generates responses like “from frame 000 to frame 200” even when the input video contains only 19 frames.
Qwen2-VL-72B. A larger-scale Vid-LLM, Qwen2-VL-72B [55], as shown in Figure 8(b), also exhibits significant hallucination issues. The model produces illogical outputs, such as incomplete sentences like “The given query happens in344-,” further emphasizing its struggle with coherent and accurate temporal reasoning.
LLaVA-Video-7B. Figure 9(a) displays the distribution of the top 10 most common time intervals predicted by LLaVA-Video-7B [71]. The model frequently outputs very short segments, such as $[1,3]$, $[2,4]$, and $[2,3]$, which together account for over 50% of predictions. This behavior suggests a significant bias in the model toward producing overly simplistic temporal spans.
LLaVA-OneVision-7B. Figure 9(b) presents the 10 most frequently predicted intervals of LLaVA-OneVision-7B [29]. The outputs are dominated by segments like $[10,12]$ and $[1,3]$, which together account for over 50% of all results. The repetitive predictions indicate a substantial hallucination problem limiting temporal reasoning capability.
7.2 VTG-Tuned Vid-LLMs
VTimeLLM. Figure 9(c) shows the predictions on the Charades-STA dataset of VTimeLLM [22]. The model frequently predicts certain frame intervals, such as $[17,34]$, which accounts for 49.34% of predictions, and $[0,17]$, which constitutes 16.34%. This pattern suggests significant hallucination and overfitting to specific frame numbers.
TimeChat. As shown in Figure 9(d), we analyze the output of the TimeChat [47] model. It tends to produce results in multiples of 5, such as 5, 10, 15, and 20. Notably, intervals like $[0,5]$ and $[0,10]$ appear in over 42% of predictions. This indicates both hallucinations and overfitting.
Table 5: The ablation results on QVHighlights dataset.
Model QVHighlights
mAP HIT@1
LLaVA-OneVision-7B 17.2 19.9
        +NumPro $20.9_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+3.7)}}}$ $27.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+7.7)}}}$
LLaVA-Video-7B 20.7 34.8
        +NumPro $22.3_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+1.6)}}}$ $38.4_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+4.4)}}}$
Qwen2-VL-72B 21.6 37.5
        +NumPro $24.2_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+2.6)}}}$ $44.3_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+6.8)}}}$
LongVA-7B-DPO 14.2 20.4
        +FT 21.9 30.8
        +NumPro-FT $25.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+10.8)}}}$ $37.2_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+16.8)}}}$
Table 6: Ablation study with different font size of NumPro-FT.
Size Color Position Charades-STA
R@0.3 R@0.5 R@0.7 mIoU
40 Red Bottom Right 63.8 42.0 20.6 41.4
60 Red Bottom Right 56.0 37.6 20.6 40.9
8 Additional Attention Analysis
Table 7: Performance comparison between the original Qwen2-VL-7B and the “Attention Map” method, which selects the two frames with the highest attention scores as the temporal boundaries.
Method Charades-STA
R@0.3 R@0.5 R@0.7 mIoU
Qwen2-VL-7B 8.7 5.4 2.4 7.9
Attention Map 18.1 11.4 3.1 19.8
We present additional attention analysis results in Figure 10. In the examples on the left, the model produces incorrect or incomplete outputs, such as “from 2 to .”. On the right, the examples display severe hallucinations, with outputs extending beyond the video’s actual duration. Despite these issues, the attention maps in both cases consistently highlight the relevant video segments. These findings show that while Vid-LLMs identify correct video segments, they fail to output accurate temporal boundaries due to the inability to translate these segments into precise textual locations.
To further examine this challenge of Vid-LLMs quantitatively, we conduct an experiment with the Qwen2-VL-7B model on the Charades-STA dataset. In this experiment, the two frames with the highest attention scores are selected as the predicted start and end frames of the segment. Specifically, we selected the two frames with the highest attention scores as the start and end frames, treating these as the predicted segment boundaries. The results, presented in Table 7, showing that this naïve attention-based solution achieves an improvement of 11.9% in mIoU compared to direct predictions of the original model. This significant gain supports our observation that Vid-LLMs have the inherent capacity to locate relevant video segments but struggle to express temporal boundaries accurately in text.
Overall, these analyses highlight the primary bottleneck of Vid-LLMs in addressing temporal grounding tasks and emphasize the need to overcome this limitation, which is what our proposed NumPro method is designed to address.
9 Ablation Results on Highlight Detection
We present additional ablation results on the QVHighlights dataset, as shown in Table 5. The results show that our method generalizes well across various General Vid-LLMs, achieving notable improvements in both mAP and HIT@1 metrics. Specifically, fine-tuning with NumPro-FT obtains a 10.8% increase in mAP and a 16.8% increase in HIT@1, surpassing state-of-the-art results.
10 Ablation Results on NumPro-FT Designs
In this section, we present the ablation results for NumPro-FT. While a font size of 60 achieves better Number Accuracy than a size of 40 (Figure 5 of the main paper), it reduces Caption Accuracy and introduces more outliers. Table 6 further supports this finding, showing that a font size of 60 results in generally lower performance, including a 7.8% drop in R@0.3 compared to size 40. We attribute this to interference with the model’s understanding of video content.
11 Additional Visualization Cases
11.1 Dialogue
Figure 11 illustrates a real-world application of our NumPro method within the Qwen2-VL-7B model, highlighting its ability to handle complex video-based dialogue tasks. Compared to the VTG-specific models [9, 34], NumPro equipped with Vid-LLMs facilitates multi-turn dialogue that adapt to user queries in real-time. For instance, the model can track score changes across video segments, identify celebrities through advanced facial recognition, and even extract textual information via OCR. These capabilities demonstrate the enhanced contextual understanding and practical value of Vid-LLMs for video comprehension tasks. By integrating NumPro, our approach further refines the temporal grounding process, enabling more precise and interactive video analysis for real-world applications.
11.2 Moment Retrieval
Figure 12 showcases additional visualization examples highlighting the effectiveness of our method in moment retrieval tasks. Our approach demonstrates robust temporal grounding capabilities by accurately identifying event boundaries across videos of varying lengths and content. Compared to previous state-of-the-art methods, it achieves substantial performance improvements.
11.3 Highlight Detection
Highlight detection [28] focuses on identifying video segments that match a given query while also assessing their relative importance. The model generates timestamps for the relevant segments and assigns a saliency score on a scale from 1 to 5. As shown in Figure 13, our method excels in accurately predicting segment start and end times, achieving consistently high mAP values. Additionally, it demonstrates precision in saliency score assessment, highlighting its suitability for tasks requiring detailed temporal localization and importance evaluation.
12 Limitations
While NumPro and NumPro-FT have proven effective across multiple models and datasets, significantly surpassing previous state-of-the-art models, there are still some limitations:
  • Limited Dataset Scope: Current datasets for video temporal grounding (VTG) tasks are predominantly focused on short videos, typically ranging from 30 seconds to 3 minutes in duration. Expanding evaluation to include longer videos, such as hour-long recordings, is essential for testing the scalability and generalizability of our approach.
  • Potential Visual Obstruction: Although NumPro is designed to minimize its impact on video content, there are scenarios where it might obscure critical visual elements, such as details, watermarks, or logos. Future enhancements could involve dynamic adjustments to the opacity of numbers or the implementation of adaptive number positioning to avoid blocking essential visual information.
  • Frame Rate Optimization: The effect of different sampling frame rates on performance remains underexplored. This study used a fixed frame rate of 1 FPS for NumPro, which may not be universally optimal. Investigating adaptive frame rates that align with the perceptual and computational characteristics of various models could lead to further improvements in accuracy and efficiency.
(a) Qwen2-VL-7B
(b) Qwen2-VL-72B
Figure 8: Video temporal grounding results where the models exhibit serious hallucination and output incorrect results. In all cases, frames are sampled at 1 FPS.
(a) LLaVA-Video-7B
(b) LLaVA-OneVision-7B
(c) VTimeLLM
(d) TimeChat
Figure 9: Distribution of the Top 10 Most Common Time Intervals within the Charades-STA Dataset for Different Models.
Figure 10: Additional attention analysis results on general Vid-LLMs
Figure 11: An example of multi-turn dialogue using our NumPro method in the Qwen2-VL-7B model.
Figure 12: Additional visualization cases of Video Temporal Grounding task.
Figure 13: Additional visualization cases of Highlight Detection task on QVHighlights dataset.

### v3

@@
 However, they struggle to extend this visual understanding to tasks requiring precise temporal localization, known as Video Temporal Grounding (VTG).
-To address this gap, we introduce Number-Prompt (NumPro), a novel method that empowers Vid-LLMs to bridge visual comprehension with temporal grounding by adding unique numerical identifiers to each video frame.
+To address this, we introduce Number-Prompt (NumPro), a novel method that empowers Vid-LLMs to bridge visual comprehension with temporal grounding by adding unique numerical identifiers to each video frame.
 Treating a video as a sequence of numbered frame images, NumPro transforms VTG into an intuitive process: flipping through manga panels in sequence.
@@
 Furthermore, fine-tuning on a NumPro-enhanced dataset defines a new state-of-the-art for VTG, surpassing previous top-performing methods by up to 6.9% in mIoU for moment retrieval and 8.5% in mAP for highlight detection.
-The code will be available at https://github.com/yongliang-wu/NumPro.
+The code is available at https://github.com/yongliang-wu/NumPro.
 Figure 1: Effectiveness of Adding Frame Numbers for Temporal Grounding: (a) Without numbered images or frames, both humans and Vid-LLMs struggle to locate specific timestamps accurately.
 (b) Once numbered, grounding temporal cues becomes as intuitive as flipping manga, where timestamps are accessible at a glance.
+footnotetext: 11footnotemark: 1 Equal Contribution.  33footnotemark: 3 Project Leader.  44footnotemark: 4 Corresponding Author.
+22footnotemark: 2 Work done during an internship at WeChat Vision, Tencent Inc.
 1 Introduction
@@
 While recognizing such actions is feasible, translating that visual information into precise timing, i.e., a specific second or frame number, is surprisingly difficult.
-This challenge is central to the field of Video Temporal Grounding (VTG) [17, 4, 43, 47, 22].
-In the realm of Video Large Language Models (Vid-LLMs) [55, 36, 71, 29], the integration of VTG allows for fine-grained visual and temporal understanding and reasoning of videos, which is pivotal for developing end-to-end video dialogue systems.
+This challenge is central to the field of Video Temporal Grounding (VTG) [18, 4, 52, 58, 25, 36].
+In the realm of Video Large Language Models (Vid-LLMs) [54, 66, 43, 89, 35, 84] which process videos as a sequence of frame images, the integration of VTG allows for fine-grained visual and temporal understanding and reasoning of videos, which is pivotal for developing end-to-end video dialogue systems.
 Despite advances of Vid-LLMs, endowing these models with effective VTG abilities presents a unique challenge: enhancing the model’s visual recognition of an event within a video does not inherently enable it to describe when the event begins and ends using language [25, 58].
 For instance, advanced Vid-LLMs like Qwen2-VL [66], while excelling at video comprehension, can struggle with grounding specific events in time.
-When asked, e.g., to locate “when does the woman eat food” in a 10-frame video, the model can hallucinate an illogical answer like “from frame 000 to 580.” [^1] See more cases in Appendix 7.1.
+When asked, e.g., to locate “when does the woman eat food” in a 10-frame video, the model can hallucinate an illogical answer like “from frame 000 to 580.”***See more cases in Appendix 7.1.
 This limitation arises because these models are primarily trained to align visual content with language descriptions (what happens) while lacking mechanisms to directly interpret the temporal boundaries (when does it happen).
+å
 This gap in powerful Vid-LLMs leads us to think: How can we empower Vid-LLMs to extract temporal cues directly through visual recognition?
@@
 Inspired by this, we introduce Number-Prompt (NumPro), which places unique numerical identifiers on each video frame, similar to manga panel numbers.
-Current Vid-LLMs process videos by considering them as a sequence of frame images.
 With NumPro, VTG is as intuitive as flipping manga.
 As shown in Figure 1, NumPro augments each frame with a unique numerical identifier denoting its position in the temporal sequence.
-Given a language query targeting an event, Vid-LLMs retrieve relevant visual features of video frames and associate with the frame numbers overlaid.
+Given a language query targeting an event, Vid-LLMs retrieve relevant visual features of video frames and associate them with the frame numbers overlaid.
 These numerical identifiers are then directly translated into textual outputs.
-In practice, we strategically position frame numbers in the bottom-right corner, using a medium font size and distinct color.
+In practice, we strategically position frame numbers in the bottom-right corner, using a defined font size and distinct color.
 This design ensures numbers visibility without obstructing essential visual content.
-Overall, NumPro allows Vid-LLMs to “read” the video timeline, effectively converting visual recognition into a temporal narrative in language.
+Overall, NumPro allows Vid-LLMs to “read” the video timeline, effectively converting visual recognition into a temporal narrative.
 NumPro’s elegance lies in its simplicity: by subtly adding frame numbers as temporal markers into video frames, we enable Vid-LLMs to naturally correlate each frame to its temporal location in the video sequence.
@@
 Our contributions can be summarized as follows:
-  • We introduce NumPro, a novel approach that enhances Video Temporal Grounding (VTG) capabilities of Video Large Language Models (Vid-LLMs) by overlaying frame numbers onto video frames, making temporal grounding as intuitive as following numbered panels in flipping manga.
+  • We introduce NumPro, a novel approach that enhances Video Temporal Grounding (VTG) capabilities of Vid-LLMs by overlaying frame numbers onto video frames, making temporal grounding as intuitive as following numbered panels in flipping manga.
   • Through an experimental study, we find a suitable NumPro design (font size, color, and position) that ensures high detectability by the model while minimally interfering with the original video content.
   • We thoroughly evaluate NumPro on standard VTG benchmarks and metrics in both training-free and fine-tuned scenarios, demonstrating its effectiveness across various models and datasets.
+Figure 2: Attention Analysis between Video Frames and Event Query.
+Although the model accurately attends to regions of interest related to the query, it struggles to generate precise temporal boundaries in its response.
+Figure 3: Framework of Our Approach in Two Settings: (1) Training-free VTG with NumPro, where frame numbers are directly added to video frames, enabling Vid-LLMs to locate events temporally without additional training, and (2) Fine-tuned VTG with NumPro-FT, which further improves VTG performance by fine-tuning Vid-LLMs on a dataset NumPro-enhanced with no architectural modifications.
 2 Related Work
 Video Temporal Grounding with Vid-LLMs.
-Video Temporal Grounding (VTG) [43, 54, 72, 32] focuses on the precise identification of event timestamps within videos, covering tasks such as moment retrieval [17, 16, 3, 68, 53, 4, 7, 63, 14, 62, 37], dense captioning [26, 24, 64, 53, 4, 8], and highlight detection [28, 47].
+Video Temporal Grounding (VTG) [52, 65, 90, 39] focuses on the precise identification of event timestamps within videos, covering tasks such as moment retrieval [18, 17, 3, 86, 64, 4, 7, 76, 15, 75, 45, 83, 81, 82, 34], dense captioning [31, 28, 77, 64, 4, 9, 19, 85, 22, 56, 67], and highlight detection [33, 58].
 For current Video Large Language Models (Vid-LLMs) [66, 89, 35], which leverage powerful LLMs [1] for cross-modal understanding and video-based reasoning, VTG is crucial for achieving fine-grained temporal and visual comprehension, enabling end-to-end video dialogue systems with integrated temporal reasoning [25, 58, 68].
@@
 While prior works focus on enhancing the region-based visual understanding of VLMs with visual prompts, our NumPro is the first to employ simple numerical tags as visual prompts within video frames to improve the temporal grounding capability.
-Figure 2: Attention Analysis between Video Frames and Event Query.
-Although the model accurately attends to regions of interest related to the query, it struggles to generate precise temporal boundaries in its response.
-Figure 3: Framework of Our Approach in Two Settings: (1) Training-free VTG with NumPro, where frame numbers are directly added to video frames, enabling Vid-LLMs to locate events temporally without additional training, and (2) Fine-tuned VTG with NumPro-FT, which further improves VTG performance by fine-tuning Vid-LLMs on a dataset NumPro-enhanced with no architectural modifications.
 3 Number-Prompt Approach
@@
 As illustrated in Figure 2, the attention map reveals a strong correlation between the text query of an event and targeted video segments.
-It indicates that Qwen2-VL-7B can effectively focus on query-relevant frames, which is consistent with the model’s strong performance in other content-related video understanding tasks [15, 30].
-However, the model struggles to verbalize the correct temporal boundaries, and generates surprising hallucinations such as “from 200 to 599.” [^2] See more cases in Appendix 8.
-This observation underscores the need for mechanisms that can bridge the gap between spatial feature alignment and temporal reasoning with Vid-LLMs, which we aim to address with our NumPro method.
+It indicates that Qwen2-VL-7B can effectively focus on query-relevant frames, which is consistent with the model’s strong performance in other content-related video understanding tasks [16, 37, 74].
+However, the model struggles to verbalize the correct temporal boundaries, and generates surprising hallucinations such as “from 200 to 599.”See more cases in Appendix 8.
+This observation underscores the need for mechanisms that can bridge the gap between spatial feature alignment and temporal reasoning with Vid-LLMs, which we aim to address.
 3.2 NumPro and NumPro-FT
@@
 By utilizing the built-in Optical Character Recognition (OCR) capabilities of Vid-LLMs, we enable them to “read” the timeline through the frame numbers associated with visual content.
-To clarify the purpose of the added numbers to Vid-LLMs, we prepend a simple instruction to each event query: “The red numbers on each frame represent the frame number.” This approach allows Vid-LLMs to accurately identify frame-level boundaries by directly linking the frame numbers to language queries.
+To clarify the purpose of the added numbers to Vid-LLMs, we prepend a simple instruction to each event query: “The red numbers on each frame represent the frame number.” This approach allows Vid-LLMs to identify frame-level boundaries by directly linking the frame numbers to language queries.
 For improved performance, NumPro-FT fine-tunes Vid-LLMs on a NumPro-augmented dataset.
@@
 To make the design choices robust across various models and datasets, we employ CLIP-based experiments on a subset of MSCOCO dataset [42] to calculate Number Accuracy and Caption Accuracy separately.
-We use the CLIP ViT-B/32 [46, 11] model to generate visual and textual representations, as many Vid-LLMs utilize CLIP-style vision encoders [55, 12], allowing our findings to generalize well across Vid-LLMs.
+We use the CLIP ViT-B/32 [57, 12, 46, 8, 91, 30, 27] model to generate visual and textual representations, as many Vid-LLMs utilize CLIP-style vision encoders [66, 13, 44], allowing our findings to generalize well across Vid-LLMs.
 COCO image-caption pairs serve as proxies for video frames, avoiding the high costs and limited scalability of direct VTG testing.
@@
 Configurations balancing these accuracies are optimal for NumPro design.
-Figure 4: Illustration of Our NumPro Design Algorithm.We overlay different numbers onto COCO images and obtain visual and textual representations using CLIP encoders.
-For each configuration, we calculate Number/Caption Similarity and derive Number/Caption Accuracy, enabling us to identify the optimal NumPro design that balances recognizability and minimal disruption to the visual content.
+Figure 4: Our NumPro Design Algorithm.
+We overlay different numbers onto COCO images and obtain visual and textual representations using CLIP encoders.
+For each configuration, we calculate Number/Caption Similarity and derive Number/Caption Accuracy, to identify the optimal NumPro design that balances recognizability and minimal disruption to the visual content.
 Figure 5: The Impact of Different Number-Prompt Designs.
@@
 Finally, we select a font size of 40, the color red, and the bottom-right position for our final NumPro design.
-This design search allows NumPro to better harness the inherent OCR and visual-language alignment capabilities of Vid-LLMs to empower video temporal grounding.
 In practice, CLIP-based designs provide approximate rather than definitive guidance, further testing on Vid-LLMs with a VTG dataset may yield additional model-specific insights.
@@
 The LLM component utilizes LoRA with parameters $r=64$ and $\alpha=128$.
-All experiments are conducted on 8 H100 GPUs.
+All experiments are conducted on 8 H800 GPUs.
 4.2 Main Results
@@
 We conduct moment retrieval experiments on Charades-STA [18] with Qwen2-VL-7B [66] in a training-free setting.
-As shown in Table 3, the results align closely with our initial observations from COCO dataset, confirming the effectiveness of our design choices in VTG tasks.
+As shown in Table 3, the results align closely with our initial observations from the COCO dataset, confirming the effectiveness of our design choices in VTG tasks.
 Specifically, (1) Position: Consistent with our CLIP-based findings, placing the text in the center has the largest impact on performance due to overlaps, while our choice of the bottom-right performs comparably to the best position; (2) Font Size: Both very large and very small fonts yield suboptimal results, supporting our balanced selection; (3) The performance on VTG is sensitive to number color, yet the red color consistently delivers the best performance, which may attribute to its high contrast against typical backgrounds [60].
@@
 Please refer to Appendix 11 for the ablation results of NumPro-FT.
+We also try directly overlaying timestamps (e.g., “10.5s”) on frames, which show inferior performance than frame numbers (Appendix 12).
 Figure 7: Performance Comparison of Sampling Strategies for NumPro.
@@
 This dual capability allows us to harness a powerful Vid-LLM for end-to-end video understanding that can flexibly adapt to both general and temporally nuanced questions within conversational AI systems.
+Moreover, we examine NumPro on more video-QA benchmarks including MVBench [37] and VideoMME [16], and we show Vid-LLMs enhanced with NumPro achieve robust performance across a variety of downstream tasks.
+Details can be found in Appendix 9.
 Table 4: The influence of applying NumPro to general video-QA.
@@
 Besides, the minimal impact on general video-QA shows that NumPro can augment VTG while maintaining robust video understanding.
+Acknowledgement
+This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), the Fundamental Research Funds for the Central Universities (2242024k30035), and the Southeast University Big Data Computing Center.
 Supplementary Material
@@
 This indicates both hallucinations and overfitting.
-Table 5: The ablation results on QVHighlights dataset.
+Table 5: The ablation results on the QVHighlights dataset.
 Model QVHighlights
@@
 Overall, these analyses highlight the primary bottleneck of Vid-LLMs in addressing temporal grounding tasks and emphasize the need to overcome this limitation, which is what our proposed NumPro method is designed to address.
+9 Additional Video Benchmark Results
+We conducted experiments on additional video question answering (QA) benchmarks, MVBench [37] and VideoMME [16], as summarized in Table 8.
+We use 1FPS as the sampling rate, and adopt a design of red color, font size 40, and bottom right positioning for the number prompt.
+Our results demonstrate that Vid-LLMs enhanced with NumPro achieve robust performance across diverse downstream tasks.
+Notably, NumPro significantly improves the Vid-LLMs’ generalization capabilities in temporal understanding tasks, such as Scene Transition and Temporal Perception.
+These findings align with our previous results presented in Table 4 on VideoInstruct [50] in the main paper.
+Table 8: Evaluations on two video QA benchmarks: MVBench and VideoMME.
+The results demonstrate that our NumPro approach enhances Vid-LLMs’ generalization capabilities on downstream tasks involving temporal understanding.
+MVBench Video-MME
+Scene Transition State Change Overall Temp. Per. Temp. Rea. Overall
+$80.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+2.5)}}}$ $42.0_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+1.0)}}}$ $51.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+0.2)}}}$ $72.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+12.7)}}}$ $49.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+8.8)}}}$ $63.7_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+0.3)}}}$
 10 Ablation Results on Highlight Detection
@@
 We attribute this to interference with the model’s understanding of video content.
+Table 9: Comparison between overlaying timestamps with overlaying frame numbers in NumPro design.
+Dataset R@0.3 R@0.5 R@0.7 mIoU
+Charades-STA $58.4_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-2.3)}}}$ $37.8_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+1.0)}}}$ $16.6_{\hskip 1.42271pt{\color[rgb]{0.1953125,0.796875,0}\textbf{(+0.7)}}}$ $37.6_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-0.9)}}}$
+ActivityNet $31.6_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-12.6)}}}$ $18.1_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-6.3)}}}$ $10.2_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-4.2)}}}$ $23.0_{\hskip 1.42271pt{\color[rgb]{1,0,0}\textbf{(-8.3)}}}$
+12 NumPro with Accurate Timestamps
+In the main paper, we choose frame numbers because they serve as the smallest discrete units of a video and can be directly mapped to precise timestamps using the frame sampling rate.
+In this section, we compare the performance by directly overlaying actual timestamps in videos.
+However, timestamps (e.g., “10.5s”) may introduce decimals, which can increase parsing complexity for Vid-LLMs.
+In Tabel 9, we compare minute-level temporal annotations (e.g., “01:10”) with frame numbers annotations sampled with 1FPS.
+The performance on Charades-STA is comparable, while frame numbers outperformed timestamps on ActivityNet, which includes longer videos and more annotations.
+These suggest that overlaying numbers with temporal information is an effective strategy for Vid-LLMs in temporal grounding, while frame numbers offer a simpler and more scalable solution.
 13 Additional Visualization Cases